# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v3

@@ Appendix B Measure-Zero Discontinuities of Correspondence-Driven Solutions under Gradient Descent
Suppose $g(x,y)$ is a $C^{2}$ function such that it is coercive and semi-algebraic in $y$ for any $x$, $\mathcal{M}$ is gradient descent method, and $0<\eta<1/\overline{L}_{g}$. We uniformly choose $b$ from $\nu,\nu]^{m}$, where $\nu>0$ can be any constant, and perturb the initial point $y_{0}$ by setting it to $y_{0}+b$. Then, with probability one, the set of points in $\mathcal{X}\setminus\widetilde{\mathcal{X}}$ at which $y^{\text{cd}}(x)$ is discontinuous has measure zero, where $y^{\mathrm{cd}}(x)$ is defined under the perturbed function $\widetilde{g}(x,y)$.
@@
Theorem B.3 (Fubini Theorem).
If the following holds
$\int_{A\times B}|f(x,y)|d(x,y)<\infty,$
where $A$ and $B$ are both $\sigma$-finite measure spaces, $A\times B$ is the product measurable space of $A$ and $B$, and $f:A\times B\to\mathbb{R}$ is a measurable function, then
@@
Detailed proof of Theorem B.1: We may assume without loss of generality that $\mathcal{X}\setminus\widetilde{\mathcal{X}}$ is connected; otherwise, since it has at most countably many connected components, the following argument can be repeated on each component and the conclusions then follow by a countable union. Since $g(x,y)$ is coercive and semi-algebraic in $y$, the gradient descent method with step size $\eta$ generates a bounded sequence that converges to a stationary point for any initialization (Attouch et al., 2013). For any fixed $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}$, $g(x,y)$ is a Morse function; thus, its stationary points are non-degenerate and isolated. Due to coercivity, the set of stationary points is bounded and therefore finite. We denote these points by $\{p_{j}(x)\}_{j\in J}$.
The basin of attraction for each stationary point $p_{j}(x)$ corresponds to its stable manifold $\mathcal{W}^{s}(p_{j}(x))$. According to the Global Stable Manifold Theorem, the stable manifold of a local minimizer is an open $m$-dimensional submanifold, whereas the stable manifolds of non-minimizing stationary points (saddle points and local maximizers) have dimension strictly less than $m$.
Discontinuities in $y^{\text{cd}}(x)$ arise only if the initialization $y_{0}$ lies in the stable manifold of a non-minimizing stationary point. Let $J^{\prime}\subset J$ denote the indices corresponding to such non-minimizers. We define the critical set of initializations as $\mathcal{A}(x):=\bigcup_{j\in J^{\prime}}\mathcal{W}^{s}(p_{j}(x))$. We introduce the indicator function for the perturbed initialization falling into this set:
$I_{\mathcal{A}}(x,b)=\mathbb{I}\{y_{0}+b\in\mathcal{A}(x)\}.$
To apply Fubini’s theorem, we must establish that $I_{\mathcal{A}}(x,b)$ is a measurable function. We define the graph of the critical set in the product space as $\mathcal{G}=\{(x,y):y\in\mathcal{A}(x)\}$. A point $(x,y)$ belongs to $\mathcal{G}$ if and only if the gradient descent sequence initiated at $y$ under the parameter $x$ converges to some non-minimizing stationary point $p_{j}(x)$ for $j\in J^{\prime}$. Let $\Phi_{k}(x,y)$ denote the $k$-th iterate of gradient descent. Since $g$ is $C^{2}$ and $\nabla_{y}g$ is $C^{1}$, $\Phi_{k}(x,y)$ is continuous jointly in $(x,y)$. Furthermore, by the implicit function theorem, the stationary points $p_{j}(x)$ depend continuously on $x$. The convergence condition can be expressed using the Borel hierarchy:
@@
$\mathcal{G}=\bigcup_{j\in J^{\prime}}\bigcap_{m=1}^{\infty}\bigcup_{N=1}^{\infty}\bigcap_{k=N}^{\infty}\left\{(x,y)\in(\mathcal{X}\setminus\widetilde{\mathcal{X}})\times\mathbb{R}^{m}:\|\Phi_{k}(x,y)-p_{j}(x)\|<\frac{1}{m}\right\}.$
The set inside the braces is the inverse image of an open interval under a continuous function, making it an open set in the product space. Since $\mathcal{G}$ is formed by countable unions and intersections of open sets, it is a Borel set. Consequently, the set $\mathcal{S}=\{(x,b):(x,y_{0}+b)\in\mathcal{G}\}$ is also a Borel set, which implies that the indicator function $I_{\mathcal{A}}(x,b)$ is Borel measurable (and thus Lebesgue measurable).
For any fixed $x$, $\mathcal{A}(x)$ is a finite union of submanifolds with dimension less than $m$, so it has Lebesgue measure zero in $\mathbb{R}^{m}$. Therefore, integrating with respect to $b$:
@@
By Fubini’s theorem, we can swap the order of integration:
$\int_{[-\nu,\nu]^{m}}\left(\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}I_{\mathcal{A}}(x,b)\,dx\right)db=\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}\left(\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}}(x,b)\,db\right)dx=0.$ (B.1)
This implies that for almost every perturbation $b\in[-\nu,\nu]^{m}$, the integral $\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}I_{\mathcal{A}}(x,b)\,dx$ is zero. Since the integrand is non-negative, this means the set of $x$ where $y^{\text{cd}}(x)$ is discontinuous has Lebesgue measure zero.

### v4

@@ Appendix B Measure-Zero Discontinuities of Correspondence-Driven Solutions under Gradient Descent
Suppose $g(x,y)$ is a $C^{2}$ function that is coercive in $y$ for any $x$. Let $\mathcal{M}$ be the gradient descent method with step size $0<\eta<1/\overline{L}_{g}$, where $\overline{L}_{g}$ is the Lipschitz smoothness coefficient of $g(x,\cdot)$. We uniformly choose $b$ from $\nu,\nu]^{m}$, where $\nu>0$ can be any constant, and perturb the initial point $y_{0}$ by setting it to $y_{0}+b$. Then, with probability one over the choice of $b$, the set of points in $\mathcal{X}\setminus\widetilde{\mathcal{X}}$ at which $y^{\text{cd}}(x)$ is discontinuous has Lebesgue measure zero, where $y^{\mathrm{cd}}(x)$ is the solution obtained by the algorithm $\mathcal{M}$ on the function $g(x,y)$ starting from the perturbed initialization $y_{0}+b$.
@@
Theorem B.3 (Fubini-Tonelli theorem).
Let $A$ and $B$ be both $\sigma$-finite measure spaces, and $A\times B$ be their product measurable space. If $f:A\times B\to[0,\infty]$ is a non-negative measurable function, then
@@
Detailed proof of Theorem B.1: Fix an arbitrary $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}$ and consider the gradient descent sequence $\{y_{k}\}$ initiated at $y_{0}$ with a step size $0<\eta<1/\overline{L}_{g}$. We first prove that the sequence generated by the gradient descent method will converge to a stationary point. Since the gradient $\nabla_{y}g(x,\cdot)$ is $\overline{L}_{g}$-Lipschitz continuous, standard optimization theory provides the following descent inequality for consecutive iterations:
$g(x,y_{k+1})\leq g(x,y_{k})-\left(\frac{1}{\eta}-\frac{\overline{L}_{g}}{2}\right)\|y_{k+1}-y_{k}\|^{2}.$
Because $\eta<1/\overline{L}_{g}$, the coefficient $C=\left(\frac{1}{\eta}-\frac{\overline{L}_{g}}{2}\right)$ is strictly positive. This strict descent property ensures that the entire sequence $\{y_{k}\}$ is strictly confined within the initial sublevel set $\mathcal{L}(x,y_{0}):=\{y\in\mathbb{R}^{m}\mid g(x,y)\leq g(x,y_{0})\}$. By the coercivity assumption on $g(x,\cdot)$, this sublevel set is bounded, and being closed, it is compact. Summing the descent inequality from $k=0$ to infinity yields a telescoping sum:
$C\sum_{k=0}^{\infty}\|y_{k+1}-y_{k}\|^{2}\leq g(x,y_{0})-\inf_{y}g(x,y).$
Since $g(x,\cdot)$ is coercive, it is bounded from below, making the right side of the inequality finite. Thus, the general term of the series must approach zero, strictly proving that $\|y_{k+1}-y_{k}\|\to 0$, and consequently $\|\nabla_{y}g(x,y_{k})\|\to 0$ as $k\to\infty$. By assumption, $g(x,\cdot)$ is a Morse function, meaning all its stationary points are strictly isolated. Because $\mathcal{L}(x,y_{0})$ is a compact set, the intersection of these isolated stationary points with $\mathcal{L}(x,y_{0})$ must form a finite discrete set, which we denote by $\mathcal{P}(x,y_{0})$. Let $\Omega$ denote the set of all accumulation points of $\{y_{k}\}$. Since $\{y_{k}\}$ stays within the compact set $\mathcal{L}(x,y_{0})$, the Bolzano-Weierstrass theorem ensures $\Omega$ is non-empty. Furthermore, since the gradient vanishes along the sequence, any accumulation point must be a stationary point, meaning $\Omega\subseteq\mathcal{P}(x,y_{0})$. Crucially, a standard result in analysis states that if a bounded sequence satisfies $\|y_{k+1}-y_{k}\|\to 0$, its set of accumulation points $\Omega$ must be a connected set. Since $\Omega$ is a non-empty connected subset of the finite discrete set $\mathcal{P}(x,y_{0})$, it must be a singleton. Therefore, the entire sequence converges strictly to a unique stationary point.
Next, we characterize the discontinuity of $y^{\text{cd}}(x)$. The basin of attraction for each stationary point corresponds to its stable manifold. According to the Global Stable Manifold Theorem, the stable manifold of a local minimizer is an open $m$-dimensional submanifold, whereas the stable manifolds of non-minimizing stationary points (saddle points and local maximizers) have dimension strictly less than $m$. If the initialization $y_{0}$ lies in the interior of the basin of attraction of a local minimizer for parameter $x$, it is easy to see that for sufficiently small perturbations of $x$, the gradient descent sequence will still converge to the continuous continuation of that local minimizer. Thus, discontinuities in $y^{\text{cd}}(x)$ arise only if the initialization $y_{0}$ lies exactly on the boundary separating different basins of attraction, which corresponds precisely to the stable manifolds of non-minimizing stationary points. Let $\Phi_{k}(x,y)$ denote the $k$-th iterate of gradient descent initiated at $y$ under the parameter $x$. As established, for any $(x,y)\in(\mathcal{X}\setminus\widetilde{\mathcal{X}})\times\mathbb{R}^{m}$, the sequence converges strictly to a stationary point. Thus, the pointwise limit function $\Phi_{\infty}(x,y):=\lim_{k\to\infty}\Phi_{k}(x,y)$ is well-defined. Since $\Phi_{k}(x,y)$ is continuous jointly in $(x,y)$ for each finite $k$, its pointwise limit $\Phi_{\infty}(x,y)$ is a Borel measurable function. We define the critical set of initializations in the product space, $\mathcal{G}$, as the set of all pairs $(x,y)$ such that the gradient descent sequence converges to a non-minimizing stationary point. A stationary point is a non-minimizer if and only if the minimal eigenvalue of the Hessian matrix of $g(x,\cdot)$ at that point is strictly negative. Therefore, $\mathcal{G}$ can be rigorously defined as:
@@
$\mathcal{G}=\left\{(x,y)\in(\mathcal{X}\setminus\widetilde{\mathcal{X}})\times\mathbb{R}^{m}:\lambda_{\min}\Big(\nabla^{2}_{yy}g\big(x,\Phi_{\infty}(x,y)\big)\Big)<0\right\}.$
Since $g$ is a $C^{2}$ function, the minimal eigenvalue function $\lambda_{\min}(\nabla^{2}_{yy}g(x,y))$ is continuous. The composition of this continuous function with the Borel measurable function $\Phi_{\infty}(x,y)$ ensures that the map $(x,y)\mapsto\lambda_{\min}\big(\nabla^{2}_{yy}g(x,\Phi_{\infty}(x,y))\big)$ is Borel measurable. Consequently, $\mathcal{G}$ is the preimage of the open interval $(-\infty,0)$ under a measurable function, making $\mathcal{G}$ a Borel set. For a fixed $x\in\mathcal{X}\setminus\widetilde{\mathcal{X}}$, let $\mathcal{A}(x)=\{y\in\mathbb{R}^{m}:(x,y)\in\mathcal{G}\}$ be the $x$-section of $\mathcal{G}$. Since each non-minimizing stationary point has a stable manifold of dimension strictly less than $m$, $\mathcal{A}(x)$ is an at most countable union of lower-dimensional submanifolds. Because the countable union of measure-zero sets still has measure zero, $\mathcal{A}(x)$ has Lebesgue measure zero in $\mathbb{R}^{m}$. We introduce the indicator function for the perturbed initialization $y_{0}+b$ falling into this critical set:
$I_{\mathcal{A}}(x,b)=\mathbb{I}\{y_{0}+b\in\mathcal{A}(x)\}=\mathbb{I}\{(x,y_{0}+b)\in\mathcal{G}\}.$
Because $\mathcal{G}$ is a Borel set, the function $(x,b)\mapsto I_{\mathcal{A}}(x,b)$ is jointly measurable. Notice that the condition $y_{0}+b\in\mathcal{A}(x)$ is strictly equivalent to $b\in\mathcal{A}(x)-y_{0}$. Since the Lebesgue measure is translation-invariant, the shifted set $\mathcal{A}(x)-y_{0}$ also has Lebesgue measure zero in $\mathbb{R}^{m}$. Therefore, integrating the indicator function with respect to the perturbation $b$ over any domain, including $\nu,\nu]^{m}$, identically yields zero:
@@
By Fubini’s theorem, since the indicator function is non-negative and measurable, we can swap the order of integration:
$\int_{[-\nu,\nu]^{m}}\left(\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}I_{\mathcal{A}}(x,b)\,dx\right)db=\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}\left(\int_{[-\nu,\nu]^{m}}I_{\mathcal{A}}(x,b)\,db\right)dx=0.$
This implies that for almost every perturbation $b\in[-\nu,\nu]^{m}$, the integral $\int_{\mathcal{X}\setminus\widetilde{\mathcal{X}}}I_{\mathcal{A}}(x,b)\,dx$ is exactly zero. Since the integrand is non-negative, this strictly means the set of $x$ where $y^{\text{cd}}(x)$ is discontinuous has Lebesgue measure zero.